### 03 Dec 2023 v1

# A Cup Product Obstruction to Frobenius Stability

Abstract.
A countable discrete group is said to be Frobenius stable if a function from the group that is “almost multiplicative” in the point Frobenius norm topology is “close” to a genuine unitary representation in the same topology. The purpose of this note is to show that if $\Gamma$ is a finitely generated group and a non-torsion element of $H^{2}(\Gamma;\mathbb{Z})$ can be written as a cup product of two elements in $H^{1}(\Gamma;\mathbb{Z})$ then $\Gamma$ is not Frobenius stable. A nontrivial corollary is that Thompson’s Group $F$ is not Frobenius stable. Because the argument is the same we also show this result for stability in unnormalized Schatten $p$-norms for $1<p\leq\infty$.
1. Introduction
Let $\Gamma$ be a countable discrete group and let $1\leq p\leq\infty$. Let $||\cdot||_{p}$ denote the unnormalized Schatten $p$-norm, $||M||_{p}=(\operatorname{Tr}((M^{*}M)^{p/2}))^{1/p}$ for $p<\infty$ and operator norm for $p=\infty$. Then a discrete group $\Gamma$ is stable in the unnormalized Schatten $p$-norm if for all sequences of functions $\varphi_{n}:\Gamma\rightarrow U_{n}$ the condition
(1) $||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}\rightarrow 0,\,\,\forall g,h\in\Gamma$
implies there exists a sequence of group homomorphisms $\varphi_{n}:\Gamma\rightarrow U_{n}$ so that
(2) $||\rho_{n}(g)-\varphi_{n}(g)||_{p}\rightarrow 0,\,\,\forall g\in\Gamma.$
Of particular interest is the $p=2$ case, called Frobenius stability, and the $p=\infty$ case called matricial stability. We will call a sequence of functions $(\varphi_{n})$ that satisfies condition (1) an asymptotic homomorphism. If there exist homomorphisms $(\psi_{n})$ satisfying condition (2) we say that $(\varphi_{n})$ is perturbable to homomorphisms. Frobenius stability was introduced by de Chiffre, Glebsky, Lubotzky, and Thom in [4].
The purpose of this note is to show the following observation.
Theorem 1.1.
Let $\Gamma$ be a finitely generated discrete group, and let $1<p\leq\infty$. If $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})$ so that $\alpha\smile\beta\in H^{2}(\Gamma;\mathbb{Z})$ is non-torsion then $\Gamma$ is not stable in the unnormalized Schatten $p$-norm.
In the $p=\infty$ case our result follows from the methods developed by Dadarlat in [3], though it does not follow from the main result as stated there. The techniques we use here are more similar to those developed by the author in [7] and [8]. In [8] the notion of a skinny cohomology class is used as an obstruction to Frobenius stability of nilpotent groups. Here a cohomology class $[\sigma]$ is skinny with respect to a homomorphism $\alpha:\Gamma\rightarrow\mathbb{Z}$ if the restriction of $[\sigma]$ to the kernel of $\alpha$ is a coboundary. Then $\alpha\smile\beta$ is skinny with respect to both $\alpha$ and $\beta$, motivating our main result.
Corollary 1.2.
For $1<p\leq\infty$ Thompson’s group $F$ is not stable in the unnormalized Schatten $p$-norm.
Proof.
The cohomology ring of $F$ can be computed with methods in [11]; see [2] for an explicit computation in the case of $F$. From this computation, one can see that $H^{1}(F;\mathbb{Z})\cong\mathbb{Z}^{2}$ and the cup product of the two generators is non-torsion. ∎
This is an interesting contrast to the result of Fournier-Facio and Rangarajan in [6] that $F$ is uniformly stable with respect to all submultiplicative norms. Here by uniformly stable, we mean that the convergence in conditions (1) and (2) is uniform rather than pointwise.
The basic idea of the proof is that since $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ we can view the pair $(\alpha,\beta)$ as a homomorphism from $\Gamma$ to $\mathbb{Z}^{2}$. We use a classic example of an asymptotic homomorphism of $\mathbb{Z}^{2}$ due to Voiculescu in [12] and pull it back by $(\alpha,\beta)$ to get an asymptotic homomorphism of $\Gamma$. We show that this is a projective representation [^1] Meaning a map from $\Gamma$ to unitaries whose failure to be multiplicative, $\rho(gh)\rho(h)^{-1}\rho(g)^{-1}$, is scalar-valued. of $\Gamma$ (Lemma 3.1). Then we use a winding number type argument based on the nontriviality of $\alpha\smile\beta$ to show that this approximate representation is not perturbable (Lemma 3.2).
The asymptotic representation that is not perturbable can be described as follows.
Definition 1.3.
Since $H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ we may view $\alpha$ and $\beta$ as group homomorphisms from $\Gamma$ to $\mathbb{Z}$. Then define
$\rho_{n}(g)=u_{n}^{\alpha(g)}v_{n}^{\beta(g)}$
where $u_{n}$ and $v_{n}$ are the $n\times n$ Voiculescu unitaries
$u_{n}=\begin{bmatrix}0&0&\cdots&0&0&1\\ 1&0&\cdots&0&0&0\\ 0&1&\cdots&0&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&0&0\\ 0&0&\cdots&0&1&0\end{bmatrix}\mbox{ and }v_{n}=\begin{bmatrix}\exp\left(\frac{2\pi i}{n}\right)&0&0&\cdots&0\\ 0&\exp\left(\frac{4\pi i}{n}\right)&0&\cdots&0\\ 0&0&\exp\left(\frac{6\pi i}{n}\right)&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}.$
2. Notation
There are many ways to characterize group homology and cohomology; we will give a concrete description of 1-cohomology, 2-cohomology, and 2-homology here. We will only use homology and cohomology with coefficients in $\mathbb{Z}$ and the trivial action in this paper. For more about this construction see [1, Chapter II.3].
As stated in the introduction, $H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ and we can take this to be the definition.
Definition 2.1.
We define a 2-cocycle to be a function $\sigma$ from $\Gamma^{2}$ to $\mathbb{Z}$ satisfying the following equation
$\sigma(g,h)-\sigma(g,hk)+\sigma(gh,k)-\sigma(h,k)=0.$
A 2-coboundary is a function that can be written in the form
$\sigma(g,h)=\gamma(g)-\gamma(gh)+\gamma(h)$
for some function $\gamma:\Gamma\rightarrow\mathbb{Z}$. Every 2-coboundary is a 2-cocycle and $H^{2}(\Gamma;\mathbb{Z})$ is defined to be the group of 2-cocycles, mod the subgroup of 2-coboundaries. The group operation is pointwise addition.
Definition 2.2.
Call $C_{k}(\Gamma)$ to be formal linear combinations of elements of $\Gamma^{k}$. We write a typical element of $C_{2}(\Gamma)$ as
$\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
with $a_{j},b_{j}\in\Gamma$ and $x_{j}\in\mathbb{Z}$. Define $\partial_{2}:C_{2}(\Gamma)\rightarrow C_{1}(\Gamma)$ to by the equation
$\partial_{2}[a|b]=[a]-[ab]+[b]$
and $\partial_{3}:C_{3}(\Gamma)\rightarrow C_{2}(\Gamma)$ by
$\partial_{3}[a|b|c]=[a|b]-[a|bc]+[ab|c]-[b|c].$
Then $H_{2}(\Gamma;\mathbb{Z}):=\ker(\partial_{2})/\operatorname{im}(\partial_{3})$. An element of $\ker(\partial_{2})$ is referred to as a 2-cycle and an element in $\operatorname{im}(\partial_{3})$ is referred to as a 2-boundary.
Definition 2.3.
The Kronecker pairing between 2-homology and 2-cohomology is a bilinear map from $H^{2}(\Gamma;\mathbb{Z})\times H_{2}(\Gamma;\mathbb{Z})$ to $\mathbb{Z}$ defined by the formula
$\left\langle\sigma,\sum_{j=1}^{N}x_{n}[a_{j}|b_{j}]\right\rangle:=\sum_{j=1}^{N}x_{j}\sigma(a_{j},b_{j})$
where $\sigma$ is a cocycle, and $\sum_{j=1}^{N}x_{n}[a_{j}|b_{j}]$ is a cycle. The value does not depend on either choice of representative.
Definition 2.4.
[…]
$\sigma(g,h)=\alpha(g)\beta(h).$
For a more general definition, and more information see [1, Chapter V.3].
Definition 2.5 ([7] Definition 3.3).
If
$c=\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]\in C_{2}(\Gamma)$
and $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\mathbb{C})$ so that for all $j\in\{1,\ldots,N\}$
$||\rho(a_{j}b_{j})\rho(a_{j})^{-1}\rho(b_{j})^{-1}-\operatorname{id}_{\mathbb{C}^{n}}||<1$
we define
$\langle\rho,c\rangle=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}(\log(\rho(a_{j}b_{j})\rho(b_{j})^{-1}\rho(a_{j})^{-1}))$
where $\log$ is defined as a power series centered at 1.
If $\partial c=0$ we have that $\langle\rho,c\rangle\in\mathbb{Z}$ [7, Proposition 3.4]. The version of the “winding number argument” we are using is as follows.
Theorem 2.6.
[7, Theorem 3.7] If $\rho_{0}$ is a (not necessarily unitary) representation of $\Gamma$, $\rho_{1}$ is a function from $\Gamma$ to $U(n)$,
$c=\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
is a 2-cycle on $\Gamma$, and
$||\rho_{1}(g)-\rho_{0}(g)||<\frac{1}{24}$
for all $g\in\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ then $\rho_{1}$ is multiplicative enough for $\langle\rho_{1},c\rangle$ to be defined and $\langle\rho_{1},c\rangle=0$.
This argument has its roots in the “winding number argument” discovered by Kazhdan [9], and later independently by Exel and Loring [5].
3. Proofs
Lemma 3.1.
Define $\rho_{n}$ as in Definition 1.3. It obeys the following identity
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}=\exp\left(-\frac{2\pi i}{n}\beta(g)\alpha(h)\right)\operatorname{id}_{\mathbb{C}^{n}}.$
Proof.
Let $\omega=\exp\left(\frac{2\pi i}{n}\right)$ and note the four identities $v_{n}u_{n}=\omega u_{n}v_{n}$, $v_{n}^{-1}u_{n}=\omega^{-1}u_{n}v_{n}^{-1}$, $v_{n}u_{n}^{-1}=\omega^{-1}u_{n}^{-1}v_{n}$, and $v_{n}^{-1}u_{n}^{-1}=\omega u_{n}^{-1}v_{n}^{-1}$. From these four identities, it follows that $v_{n}^{x}u_{n}^{y}=\omega^{xy}u_{n}^{x}v_{n}^{y}$, for all $x,y\in\mathbb{Z}$. We compute
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}$ $=u_{n}^{\alpha(g)+\alpha(h)}v_{n}^{\beta(g)+\beta(h)}v_{n}^{-\beta(h)}u_{n}^{-\alpha(h)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=u_{n}^{\alpha(g)+\alpha(h)}v_{n}^{\beta(g)}u_{n}^{-\alpha(h)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=\omega^{-\beta(g)\alpha(h)}u_{n}^{\alpha(g)+\alpha(h)}u_{n}^{-\alpha(h)}v_{n}^{\beta(g)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=\omega^{-\beta(g)\alpha(h)}\operatorname{id}_{\mathbb{C}^{n}}.$
Lemma 3.2.
Let $\Gamma$ be a countable discrete group and let $[\sigma]\in H^{2}(\Gamma;\mathbb{Z})$ be a cohomology class with cocycle representative $\sigma$. Suppose that $\rho_{n}$ is a sequence of functions from $\Gamma$ to $k_{n}\times k_{n}$ unitaries so that
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}=\exp\left(\frac{2\pi i}{n}\sigma(g,h)\right)\operatorname{id}_{\mathbb{C}^{n}}.$
Then
$||\rho_{n}(g)\rho_{n}(h)-\rho_{n}(gh)||\leq 2\pi|\sigma(g,h)|/n;$
in particular, $\rho_{n}$ is asymptotically multiplicative in operator norm.
Now suppose that $[\sigma]$ pairs nontrivially with some homology class. Then $\rho_{n}$ is not perturbable to homomorphisms in operator norm.
Proof.
This follows the proof of [7, Theorem 3.20].
The first part follows from the fact that
$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||=||\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}-\operatorname{id}_{\mathbb{C}^{k_{n}}}||.$
Now we will show that for large enough $n$, $\rho_{n}$ is not close to any genuine representation of $\Gamma$ in operator norm on a particular finite subset of $\Gamma$. There is some 2-cycle $c\in C_{2}(\Gamma)$ written
$c=\sum_{i=1}^{N}x_{i}[a_{i}|b_{i}].$
so that
$\langle\sigma,c\rangle\neq 0.$
Then we compute that
$\langle\rho_{n},c\rangle$ $=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}(\log(\rho_{n}(a_{j}b_{j})\rho_{n}(b_{j})^{-1}\rho_{n}(a_{j})^{-1}))$
$=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}\left(\log\left(\exp\left(\frac{2\pi i}{n}\sigma(a_{j},b_{j})\right)\operatorname{id}_{\mathbb{C}^{k_{n}}}\right)\right)$
$=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\frac{2\pi i}{n}\sigma(a_{j},b_{j})\operatorname{Tr}(\operatorname{id}_{\mathbb{C}^{k_{n}}})$
$=\langle\sigma,c\rangle\frac{k_{n}}{n}$
$\neq 0.$
By Theorem 2.6 it follows that $\rho_{n}$ cannot be within $\frac{1}{24}$ of a genuine representation on the boundary set $\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ and thus cannot be perturbed to a genuine representation.
Now we are ready to prove Theorem 1.1.
Proof.
Note that by the definition of the cup product $\sigma(g,h)=-\beta(g)\alpha(h)$ is a cocycle representative of the cohomology class $-\beta\smile\alpha=\alpha\smile\beta$.
First, we show asymptotic multiplicativity. In operator norm, this follows directly from Lemma 3.1 and Lemma 3.2. For the $p<\infty$ case, we use the same lemmas to show
$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}$ $\leq||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||\cdot n^{1/p}$
$\leq 2\pi|\alpha(g)\beta(h)|\cdot n^{1/p-1}.$
This goes to zero for $p>1$.
Since $\alpha\smile\beta$ is non-torsion it must also pair nontrivially with a 2-homology class; to see this note that from the universal coefficient theorem [10, Theorem 53.1] we have a short exact sequence
[…]
Since $\Gamma$ is finitely generated $H_{1}(\Gamma;\mathbb{Z})\cong\Gamma/[\Gamma,\Gamma]$ [1, page 36] is finitely generated as well. Thus $\operatorname{Ext}(H_{1}(\Gamma;\mathbb{Z}),\mathbb{Z}))$ can be show to be torsion from [10, Theorem 52.3] and the table on [10] page 331.
Now it follows from Lemma 3.1 and Lemma 3.2 that $\rho_{n}$ is not perturbable in operator norm to a sequence of genuine representations. Since operator norm is smaller than any other unnormalized Schatten $p$-norm it follows that $\rho_{n}$ cannot be close to a genuine representation in any of these norms either. ∎
Remark 3.3.
Note that the finitely generated condition in Theorem 1.1 can be dropped if we require that $\alpha\smile\beta$ pairs nontrivially with a 2-homology class.
Acknowledgments
I would like to thank Eli Bashwinger for answering a question I had about the cohomology of Thompson and Thompson-like groups; his answer inspired me to make this note. I would also like to thank the Purdue math department for supporting me with the Ross-Lynn grant during the 2023-2024 academic year.

### 07 Feb 2024 v3

@@
 Abstract.
-A countable discrete group is said to be Frobenius stable if a function from the group that is “almost multiplicative” in the point Frobenius norm topology is “close” to a genuine unitary representation in the same topology.
-The purpose of this note is to show that if $\Gamma$ is a finitely generated group and a non-torsion element of $H^{2}(\Gamma;\mathbb{Z})$ can be written as a cup product of two elements in $H^{1}(\Gamma;\mathbb{Z})$ then $\Gamma$ is not Frobenius stable.
-A nontrivial corollary is that Thompson’s Group $F$ is not Frobenius stable.
-Because the argument is the same we also show this result for stability in unnormalized Schatten $p$-norms for $1<p\leq\infty$.
+A countable discrete group $\Gamma$ is said to be Frobenius stable if a function from the group that is “almost multiplicative” in the point Frobenius norm topology is “close” to a genuine unitary representation in the same topology.
+The purpose of this paper is to show that if $\Gamma$ is finitely generated and a non-torsion element of $H^{2}(\Gamma;\mathbb{Z})$ can be written as a cup product of two elements in $H^{1}(\Gamma;\mathbb{Z})$ then $\Gamma$ is not Frobenius stable.
+In general, 2-cohomology does not obstruct Frobenius stability.
+Some examples are discussed, including Thompson’s group $F$ and Houghton’s group $H_{3}$.
+The argument is sufficiently general to show that the same condition implies non-stability in unnormalized Schatten $p$-norms for $1<p\leq\infty$.
 1. Introduction
-Let $\Gamma$ be a countable discrete group and let $1\leq p\leq\infty$.
-Let $||\cdot||_{p}$ denote the unnormalized Schatten $p$-norm, $||M||_{p}=(\operatorname{Tr}((M^{*}M)^{p/2}))^{1/p}$ for $p<\infty$ and operator norm for $p=\infty$.
-Then a discrete group $\Gamma$ is stable in the unnormalized Schatten $p$-norm if for all sequences of functions $\varphi_{n}:\Gamma\rightarrow U_{n}$ the condition
-(1) $||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}\rightarrow 0,\,\,\forall g,h\in\Gamma$
-implies there exists a sequence of group homomorphisms $\varphi_{n}:\Gamma\rightarrow U_{n}$ so that
-(2) $||\rho_{n}(g)-\varphi_{n}(g)||_{p}\rightarrow 0,\,\,\forall g\in\Gamma.$
+Let $1\leq p\leq\infty$ and let $||\cdot||_{p}$ denote the unnormalized Schatten $p$-norm on the space of $k\times k$ complex matrices, $||M||_{p}=(\operatorname{Tr}((M^{*}M)^{p/2}))^{1/p}$ for $p<\infty$ and operator norm for $p=\infty$.
+Then a countable discrete group $\Gamma$ is stable in the unnormalized Schatten $p$-norm if for all sequences of functions $\varphi_{n}$ from $\Gamma$ to the complex $k_{n}\times k_{n}$ unitary group $U_{k_{n}}$ the condition
+(1) $||\varphi_{n}(gh)-\varphi_{n}(g)\varphi_{n}(h)||_{p}\rightarrow 0,\,\,\forall g,h\in\Gamma$
+implies there exists a sequence of group homomorphisms $\psi_{n}:\Gamma\rightarrow U_{k_{n}}$ so that
+(2) $||\psi_{n}(g)-\varphi_{n}(g)||_{p}\rightarrow 0,\,\,\forall g\in\Gamma.$
 Of particular interest is the $p=2$ case, called Frobenius stability, and the $p=\infty$ case called matricial stability.
@@
 Frobenius stability was introduced by de Chiffre, Glebsky, Lubotzky, and Thom in [8].
-The purpose of this note is to show the following observation.
+Stability of a finitely presented group is equivalent to a notion of stability of the presentation of that group; this notion was shown to be independent of the presentation by Arzhantseva and Păunescu in [1].
+The goal of this paper is to show the following.
 Theorem 1.1.
 Let $\Gamma$ be a finitely generated discrete group, and let $1<p\leq\infty$.
-If $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})$ so that $\alpha\smile\beta\in H^{2}(\Gamma;\mathbb{Z})$ is non-torsion then $\Gamma$ is not stable in the unnormalized Schatten $p$-norm.
+If there are $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})$ so that $\alpha\smile\beta\in H^{2}(\Gamma;\mathbb{Z})$ is non-torsion then $\Gamma$ is not stable in the unnormalized Schatten $p$-norm.
+In general nonzero second Betti number does not obstruct Frobenius stability; in [2] Bader, Lubotzky, Sauer, and Weinberger show that $\operatorname{Sp}_{2n+2}(\mathbb{Z})$ is Frobenius stable despite having a nonzero second Betti number.
 In the $p=\infty$ case our result follows from the methods developed by Dadarlat in [7], though it does not follow from the main result as stated there.
@@
 Then $\alpha\smile\beta$ is skinny with respect to both $\alpha$ and $\beta$, motivating our main result.
-Corollary 1.2.
-For $1<p\leq\infty$ Thompson’s group $F$ is not stable in the unnormalized Schatten $p$-norm.
-Proof.
-The cohomology ring of $F$ can be computed with methods in [11]; see [2] for an explicit computation in the case of $F$.
-From this computation, one can see that $H^{1}(F;\mathbb{Z})\cong\mathbb{Z}^{2}$ and the cup product of the two generators is non-torsion.
-∎
-This is an interesting contrast to the result of Fournier-Facio and Rangarajan in [6] that $F$ is uniformly stable with respect to all submultiplicative norms.
-Here by uniformly stable, we mean that the convergence in conditions (1) and (2) is uniform rather than pointwise.
 The basic idea of the proof is that since $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ we can view the pair $(\alpha,\beta)$ as a homomorphism from $\Gamma$ to $\mathbb{Z}^{2}$.
@@
 of $\Gamma$ (Lemma 3.1).
-Then we use a winding number type argument based on the nontriviality of $\alpha\smile\beta$ to show that this approximate representation is not perturbable (Lemma 3.2).
-The asymptotic representation that is not perturbable can be described as follows.
+Then we use a winding number type argument based on the nontriviality of $\alpha\smile\beta$ to show that this asymptotic representation is not perturbable (Lemma 3.2).
+In Section 4 we go over examples of groups that Theorem 1.1 applies to.
+We show that Thompson’s group $F$ and Houghton’s groups $H_{n}$ for $n\geq 3$ among other examples are not stable in the unnormalized Schatten $p$-norm for $p>1$.
+In some cases, there is a more direct elementary argument that the group is not stable because the map $(\alpha,\beta)$ from the group to $\mathbb{Z}^{2}$ splits; this is the case for Thompson’s group $F$ in particular.
+In many cases the splittings are not obvious, so the main result is still useful in identifying the groups as non-stable.
+An asymptotic representation that is not perturbable can be described as follows.
 Definition 1.2.
@@
 We will only use homology and cohomology with coefficients in $\mathbb{Z}$ and the trivial action in this paper.
-For more about this construction see [1, Chapter II.3].
+For more about this construction see [4, Chapter II.3].
 As stated in the introduction, $H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ and we can take this to be the definition.
@@
 Definition 2.2.
-Call $C_{k}(\Gamma)$ to be formal linear combinations of elements of $\Gamma^{k}$. We write a typical element of $C_{2}(\Gamma)$ as
+Define $C_{k}(\Gamma)$ to be formal linear combinations of elements of $\Gamma^{k}$. We write a typical element of $C_{2}(\Gamma)$ as
 $\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
@@
 $\sigma(g,h)=\alpha(g)\beta(h).$
-For a more general definition, and more information see [1, Chapter V.3].
+For a more general definition, and more information see [4, Chapter V.3].
+Proposition 2.5.
+Let $q:\Gamma\rightarrow\Lambda$ be a group homomorphism.
+Then there is a map $q^{*}:H^{*}(\Lambda;\mathbb{Z})\rightarrow H^{*}(\Gamma;\mathbb{Z})$.
+Moreover $q^{*}(\alpha\smile\beta)=q^{*}(\alpha)\smile q^{*}(\beta)$.
 Definition 2.6 ([12] Definition 3.3).
@@
 and $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\mathbb{C})$ so that for all $j\in\{1,\ldots,N\}$
-$||\rho(a_{j}b_{j})\rho(a_{j})^{-1}\rho(b_{j})^{-1}-\operatorname{id}_{\mathbb{C}^{n}}||<1$
+$||\rho(a_{j}b_{j})\rho(a_{j})^{-1}\rho(b_{j})^{-1}-\operatorname{id}_{\mathbb{C}^{n}}||_{\infty}<1$
 we define
@@
 where $\log$ is defined as a power series centered at 1.
-If $\partial c=0$ we have that $\langle\rho,c\rangle\in\mathbb{Z}$ [7, Proposition 3.4]. The version of the “winding number argument” we are using is as follows.
+If $\partial c=0$ we have that $\langle\rho,c\rangle\in\mathbb{Z}$ [12, Proposition 3.4]. The version of the “winding number argument” we are using is as follows.
 Theorem 2.7.
-[7, Theorem 3.7] If $\rho_{0}$ is a (not necessarily unitary) representation of $\Gamma$, $\rho_{1}$ is a function from $\Gamma$ to $U(n)$,
+[12, Theorem 3.7] If $\rho_{0}$ is a (not necessarily unitary) representation of $\Gamma$, $\rho_{1}$ is a function from $\Gamma$ to $U(n)$,
 $c=\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
 is a 2-cycle on $\Gamma$, and
-$||\rho_{1}(g)-\rho_{0}(g)||<\frac{1}{24}$
+$||\rho_{1}(g)-\rho_{0}(g)||_{\infty}<\frac{1}{24}$
 for all $g\in\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ then $\rho_{1}$ is multiplicative enough for $\langle\rho_{1},c\rangle$ to be defined and $\langle\rho_{1},c\rangle=0$.
@@
 Then
-$||\rho_{n}(g)\rho_{n}(h)-\rho_{n}(gh)||\leq 2\pi|\sigma(g,h)|/n;$
+$||\rho_{n}(g)\rho_{n}(h)-\rho_{n}(gh)||_{\infty}\leq 2\pi|\sigma(g,h)|/n;$
 in particular, $\rho_{n}$ is asymptotically multiplicative in operator norm.
@@
 Proof.
-This follows the proof of [7, Theorem 3.20].
+This follows the proof of [12, Theorem 3.20].
 The first part follows from the fact that
-$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||=||\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}-\operatorname{id}_{\mathbb{C}^{k_{n}}}||.$
+$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{\infty}=||\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}-\operatorname{id}_{\mathbb{C}^{k_{n}}}||_{\infty}.$
 Now we will show that for large enough $n$, $\rho_{n}$ is not close to any genuine representation of $\Gamma$ in operator norm on a particular finite subset of $\Gamma$.
@@
 $\neq 0.$
-By Theorem 2.6 it follows that $\rho_{n}$ cannot be within $\frac{1}{24}$ of a genuine representation on the boundary set $\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ and thus cannot be perturbed to a genuine representation.
+By Theorem 2.7 it follows that $\rho_{n}$ cannot be within $\frac{1}{24}$ of a genuine representation on the set $\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ and thus cannot be perturbed to a genuine representation.
 Now we are ready to prove Theorem 1.1.
 Proof.
-Note that by the definition of the cup product $\sigma(g,h)=-\beta(g)\alpha(h)$ is a cocycle representative of the cohomology class $-\beta\smile\alpha=\alpha\smile\beta$.
+Note that by the definition of the cup product $\sigma(g,h)=-\beta(g)\alpha(h)$ is a cocycle representative of the cohomology class $-\beta\smile\alpha=\alpha\smile\beta$ [4, V.3 equation 3.6].
 First, we show asymptotic multiplicativity. In operator norm, this follows directly from Lemma 3.1 and Lemma 3.2. For the $p<\infty$ case, we use the same lemmas to show
-$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}$ $\leq||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||\cdot n^{1/p}$
+$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}$ $\leq||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{\infty}\cdot n^{1/p}$
 $\leq 2\pi|\alpha(g)\beta(h)|\cdot n^{1/p-1}.$
 This goes to zero for $p>1$.
-Since $\alpha\smile\beta$ is non-torsion it must also pair nontrivially with a 2-homology class; to see this note that from the universal coefficient theorem [10, Theorem 53.1] we have a short exact sequence
+Since $\alpha\smile\beta$ is non-torsion it must also pair nontrivially with a 2-homology class; to see this note that from the universal coefficient theorem [18, Theorem 53.1] we have a short exact sequence
 […]
-Since $\Gamma$ is finitely generated $H_{1}(\Gamma;\mathbb{Z})\cong\Gamma/[\Gamma,\Gamma]$ [1, page 36] is finitely generated as well.
-Thus $\operatorname{Ext}(H_{1}(\Gamma;\mathbb{Z}),\mathbb{Z}))$ can be show to be torsion from [10, Theorem 52.3] and the table on [10] page 331.
+Since $\Gamma$ is finitely generated $H_{1}(\Gamma;\mathbb{Z})\cong\Gamma/[\Gamma,\Gamma]$ [4, page 36] is finitely generated as well.
+Thus $\operatorname{Ext}(H_{1}(\Gamma;\mathbb{Z}),\mathbb{Z}))$ can be show to be torsion from [18, Theorem 52.3] and the table on [18] page 331.
 Now it follows from Lemma 3.1 and Lemma 3.2 that $\rho_{n}$ is not perturbable in operator norm to a sequence of genuine representations.
-Since operator norm is smaller than any other unnormalized Schatten $p$-norm it follows that $\rho_{n}$ cannot be close to a genuine representation in any of these norms either.
+Since the operator norm is smaller than any other unnormalized Schatten $p$-norm it follows that $\rho_{n}$ cannot be close to a genuine representation in any of these norms either.
 ∎
@@
 Note that the finitely generated condition in Theorem 1.1 can be dropped if we require that $\alpha\smile\beta$ pairs nontrivially with a 2-homology class.
+4. Examples
+Note that our result relies on the existence of a homomorphism $(\alpha,\beta)$ from $\Gamma$ to $\mathbb{Z}^{2}$.
+Because $\alpha\smile\beta=-\beta\smile\alpha$ [4, V.3 equation 3.6] our assumption that $\alpha\smile\beta$ is non-torsion implies that $\alpha$ and $\beta$ must be linearly independent.
+If the map $(\alpha,\beta)$ is a surjection that splits, our theorem applies, but in this case, the non-stability can be proven with a simpler argument.
+Since the argument is more general we will explain a more general context.
+Definition 4.1 ([8]).
+If $\{(G_{k},d_{k})\}$ is a family of groups with bi-invariant metrics, $(G_{k},d_{k})$-stability is defined analogously go stability in the unnormalized Schatten $p$-norm but $U_{k_{n}}$ is replaced by the family $G_{k_{n}}$ and the convergence in both conditions is replaced by convergence in the metric $d_{k_{n}}$.
+Uniform $(G_{k},d_{k})$-stability is defined analogously, but with uniform convergence instead of pointwise convergence.
+If the map $(\alpha,\beta):\Gamma\rightarrow\mathbb{Z}^{2}$ splits then the non-stability of $\Gamma$ can be proved with the following lemma pointed out to the author by Francesco Fournier-Facio.
+Lemma 4.2.
+Suppose that $\Lambda\rtimes\Upsilon$ is (uniformly) $(G_{k},d_{k})$-stable. Then $\Upsilon$ is (uniformly) $(G_{k},d_{k})$-stable as well.
+Proof.
+Call $\Gamma=\Lambda\rtimes\Upsilon$ and let $r:\Gamma\rightarrow\Upsilon$ and $s:\Upsilon\rightarrow\Gamma$ be the obvious maps.
+Let $\rho_{n}$ be a sequence of (uniformly) asymptotically multiplicative maps from $\Upsilon$ to $G_{k_{n}}$.
+Then $\rho_{n}\circ r$ are (uniformly) asymptotically multiplicative as well.
+Thus there is a sequence of representations $\pi_{n}:\Gamma\rightarrow G_{k_{n}}$ be a sequence of representations that (uniformly) approximate $\rho_{n}\circ r$.
+Then $\pi_{n}\circ s$ (uniformly) approximates $\rho_{n}\circ r\circ s=\rho_{n}$.
+∎
+Many examples of groups that satisfy the conditions of Theorem 1.1 have a split homomorphism to $\mathbb{Z}^{2}$, but not all; the easiest counterexample is higher genus surface groups.
+These satisfy the conditions of the theorem, but they are hyperbolic so $\mathbb{Z}^{2}$ cannot be embedded as a subgroup in them by [3, Corollary III.$\Gamma$ 3.10].
+They are already known not to be stable in unnormalized Schatten $p$-norm for $1<p\leq\infty$; it follows from the computation in the proof of Theorem 2 in [16].
+In our examples, we will prove that groups are not stable using Theorem 1.1, then use Lemma 4.2 to provide an alternate proof if possible.
+Often the splittings are not obvious, so Theorem 1.1 is still useful in identifying these groups as non-stable; often the proof that the group satisfies the cohomological condition is shorter than the proof that the map splits.
+Many interesting examples come from extensions of a group that fits the conditions of Theorem 1.1 by a locally finite group, a class of groups that was suggested to the author by Francesco Fournier-Facio.
+Lemma 4.3.
+Suppose $\Lambda\trianglelefteq\Gamma$ and let $q:\Gamma\rightarrow\Gamma/\Lambda$ be the quotient map.
+Suppose that $H^{1}(\Lambda;\mathbb{Z})=\{0\}$.
+Let $\alpha,\beta\in H^{2}(\Gamma/\Lambda;\mathbb{Z})$.
+If $\alpha\smile\beta$ is non-torsion it follows that $q^{*}(\alpha)\smile q^{*}(\beta)$ is non-torsion as well.
+Proof.
+From the exact sequence in [14] Theorem III.2 we see that the map $q^{*}:H^{2}(\Gamma/\Lambda;\mathbb{Z})\rightarrow H^{2}(\Gamma;\mathbb{Z})$ is an injection.
+Because $q^{*}$ preserves cup product we must have that $q^{*}(\alpha)\smile q^{*}(\beta)=q^{*}(\alpha\smile\beta)$ is non-torsion.
+∎
+Remark 4.4.
+Note that since $H^{1}(\Lambda;\mathbb{Z})\cong\operatorname{Hom}(\Lambda;\mathbb{Z})$ it follows that if all elements of $\Lambda$ are torsion then $H^{1}(\Lambda;\mathbb{Z})=\{0\}$.
+Thus if we have an extension
+[…]
+where $\Lambda$ has only torsion elements, $\Gamma$ is finitely generated, and $\Upsilon$ meets the conditions of Theorem 1.1 then $\Gamma$ meets the condition of Theorem 1.1 as well.
+The same argument holds if $\Lambda$ has property (T), is simple, or is any other group without 1-cohomology.
+Definition 4.5 ([15]).
+The Houghton groups $H_{n}$ for $n\geq 3$ are defined as follows. Let $X_{n}=\{1,\ldots,n\}\times\mathbb{N}$. For $k\in\{2,\ldots,n\}$ define
+$g_{k}(x,y)=\begin{cases}(x,y+1)&x=1;\\ (1,0)&(x,y)=(k,0);\\ (x,y-1)&x=k\mbox{ and }y\neq 0;\\ (x,y)&\mbox{otherwise.}\end{cases}$
+Then $H_{n}$ is the subgroup of permutations of $X_{n}$ spanned by $g_{2},\ldots,g_{n}$.
+Corollary 4.6.
+For $n\geq 3$, and $1<p\leq\infty$ the Houghton group $H_{n}$ is not stable in the unnormalized Schatten $p$-norm.
+Proof.
+It follows from an argument originally due to [21] (see [17] for a full explanation) that there is an extension
+[…]
+so by Remark 4.4 and Theorem 1.1 it follows that $H_{n}$ is not stable in the Schatten $p$-norm. ∎
+For $n\geq 4$ one can find a way to show this with Lemma 4.2 instead of Theorem 1.1.
+If $g_{2},\ldots,g_{n}$ are the generators above.
+Note that for $i$, $j$, $k$ distinct the elements $g_{i}^{-1}g_{j}$ and $g_{k}$ commute with each other.
+So the map from $H_{n}\rightarrow\mathbb{Z}^{2}$ defined by $g_{j}\mapsto e_{1}$, $g_{k}\mapsto e_{2}$, and $g_{\ell}\mapsto 0$ for $\ell\neq j,k$ has a splitting determined by $e_{1}\mapsto g_{i}^{-1}g_{j}$ and $e_{2}\mapsto g_{k}$.
+The author does not see an obvious splitting in the $n=3$ case.
+Corollary 4.7.
+For $1<p\leq\infty$ Thompson’s group $F$ is not stable in the unnormalized Schatten $p$-norm.
+Proof.
+The cohomology ring of $F$ can be computed with methods in [19]; see [5] for an explicit computation in the case of $F$.
+From this computation, one can see that $H^{1}(F;\mathbb{Z})\cong\mathbb{Z}^{2}$ and the cup product of the two generators is non-torsion.
+∎
+A similar question of whether or not $F$ is permutation stable was raised by Arzhantseva and Păunescu in [1], where they point out that if $F$ were permutation stable it would imply that it is not sofic, which would demonstrate both the existence of a non-sofic group and that $F$ is non-amenable, both of which are famous unsolved problems.
+Showing that $F$ is stable in the normalized Hilbert-Schmidt norm would also be very interesting because it would show that $F$ is not hyperlinear by the same argument they give.
+Our result is an interesting contrast to the result of Fournier-Facio and Rangarajan in [10] that $F$ is uniformly stable with respect to all submultiplicative norms.
+An alternate proof of Corollary 4.7 fact (using Lemma 4.2) was pointed out to the author by Francesco Fournier-Facio.
+View $F$ as the group piecewise-linear homeomorphisms from the unit interval to itself where all slopes are powers of 2 and all non-differentiable points are dyadic rationals.
+The group operation is composition.
+Let $f,g\in F$ be any two functions so that $f|_{[0,\frac{1}{2}]}=\operatorname{id}_{[0,\frac{1}{2}]}$ and $g|_{[\frac{1}{2},1]}=\operatorname{id}_{[\frac{1}{2},1]}$, but $f^{\prime}(1)=g^{\prime}(0)=2$.
+One can define homomorphisms $\alpha,\beta:F\rightarrow\mathbb{Z}$ by $\alpha(h)=\log_{2}(h^{\prime}(1))$ and $\beta(h)=\log_{2}(h^{\prime}(0))$.
+Then sending the generators of $\mathbb{Z}^{2}$ to $f$ and $g$ provides a splitting for the map $(\alpha,\beta)$.
+Another class of examples are groups of the form $\Gamma\vee\Lambda$. Here $\Gamma\vee\Lambda$ is defined as follows.
+Definition 4.8 ([11]).
+Pick set models of $\Gamma$ and $\Lambda$ so that the identity elements of $\Gamma$ and $\Lambda$ are identified with each other, but $(\Gamma\setminus\{e\})\cap(\Lambda\setminus\{e\})=\emptyset$.
+Then let $S$ be the set $\Gamma\cup\Lambda$.
+Then define an action of $\Gamma$ on $S$ so that $\Gamma$ acts on itself by left translation while it stabilizes elements of $\Lambda\setminus\{e\}$ and defines an action of $\Lambda$ on $S$ analogously.
+Define $\Gamma\vee\Lambda$ to be the subgroup of $\operatorname{Sym}(S)$ generated by the actions of $\Gamma$ and $\Lambda$.
+One can show that if $H^{1}(\Gamma;\mathbb{Z})$ and $H^{1}(\Lambda;\mathbb{Z})$ are both nontrivial, then $\Gamma\times\Lambda$ is not stable in the unnormalized Schatten $p$-norm, for $1<p\leq\infty$, by applying Lemma 4.2 as follows.
+By the cohomological assumption, there are maps $\alpha:\Gamma\rightarrow\mathbb{Z}$ and $\beta:\Lambda\rightarrow\mathbb{Z}$, which we may take to be surjections.
+Then the map $(\alpha,\beta):\Gamma\times\Lambda\rightarrow\mathbb{Z}^{2}$ clearly splits.
+From this it follows that $\alpha\smile\beta$ is non-torsion.
+Corollary 4.9.
+Let $1<p\leq\infty$.
+If $\Gamma$ and $\Lambda$ are finitely generated groups so that the integer cohomology of each is nontrivial then $\Gamma\vee\Lambda$ is not stable in the unnormalized Schatten $p$-norm.
+Proof.
+Note that both groups must be infinite, in order to have integer cohomology. Then by [11, Proposition 1.7], for any infinite groups $\Gamma$ and $\Lambda$ there is a short exact sequence
+[…]
+where $S$ is the set defined in Definition 4.8 and $\operatorname{Alt}_{f}(S)$ is the group of finitely-supported even permutations of $S$. From Remark 4.4 the result follows. ∎
+This can also be explained with Lemma 4.2.
+While by [11, Theorem 1.8] the map from $\Gamma\times\Lambda\rightarrow\Gamma\vee\Lambda$ often does not split, the composition with the map to $\mathbb{Z}^{2}$ does.
+If $\alpha$ and $\beta$ are surjective elements in $\operatorname{Hom}(\Gamma,\mathbb{Z})$ and $\operatorname{Hom}(\Lambda,\mathbb{Z})$ respectively then let $s:\mathbb{Z}^{2}\rightarrow\Gamma\times\Lambda$ be a splitting.
+By [11, Proposition 7.1] this also induces an, $s_{\vee}$ from $\mathbb{Z}\vee\mathbb{Z}$ into $\Gamma\vee\Lambda$.
+By [11, Proposition 4.2] the map, $\pi_{\mathbb{Z}}:\mathbb{Z}\vee\mathbb{Z}\rightarrow\mathbb{Z}^{2}$ does split; call the splitting $t$.
+Then we claim that $s_{\vee}\circ t$ is a splitting of $(\alpha,\beta)\circ\pi$.
+To show this compute
+$(\alpha,\beta)\circ\pi\circ s_{\vee}\circ t$ $=(\alpha,\beta)\circ s\circ\pi_{\mathbb{Z}}\circ t$
+$=\operatorname{id}_{\mathbb{Z}^{2}}.$
+Definition 4.10 ([6]).
+Suppose $\Upsilon$ and $\Lambda$ are groups, and $I$ is a set that $\Upsilon$ acts on.
+Then a wreath-like product of $\Lambda$ by $\Upsilon$ with corresponding to the action is a group $\Gamma$ fitting into an extension
+[…]
+where $\Lambda_{i}\cong\Lambda$, and for all $g\in\Gamma$, $g\Lambda_{i}g^{-1}=\Lambda_{\pi(g)\cdot i}$.
+Corollary 4.11.
+If $\Lambda$ is a finitely generated group with trivial 1-cohomology, and $\Upsilon$ satisfies the conditions of Theorem 1.1 then so does any wreath-like product of $\Lambda$ by $\Upsilon$ with a transitive action.
+In particular, a wreath-like product of a finite group, property (T) group, or finitely generated simple group by $\mathbb{Z}^{n}$ for $n\geq 2$, with a transitive action, is not stable in the Schatten $p$-norm for $1<p\leq\infty$.
+Proof.
+Note that any nonzero homomorphism from $\bigoplus_{i\in I}\Lambda_{i}$ must pull back to a nonzero homomorphism from $\Lambda_{i}$ to $\mathbb{Z}$ for some $i$.
+It follows that $H^{1}\left(\bigoplus_{i\in I}\Lambda_{i};\mathbb{Z}\right)=\{0\}$.
+Now the result follows from the definition and Remark 4.4, if we can show that the wreath-like product is finitely generated.
+To this end let $\Lambda$, $\Gamma$, and $\Upsilon$ be as above.
+Let $g_{1},\ldots,g_{n}\in\Gamma$ so that $\pi(g_{1}),\ldots,\pi(g_{n})$ generate $\Upsilon$, pick a fixed $j\in I$, and let $h_{1},\ldots,h_{m}$ generate $\Lambda_{j}$.
+We claim that $g_{1},\ldots,g_{k},h_{1},\ldots,h_{m}$ generates $\Gamma$.
+Let $G\leq\Gamma$ be the subgroup generated by $g_{1},\ldots,g_{k}$, and note that by definition $\pi(G)=\Upsilon$.
+Now note that for all $i\in I$ there is some $y\in\Upsilon$ so that $y\cdot j=i$.
+Then letting $g\in G$ so that $\pi(g)=y$ we have that $g\Lambda_{j}g^{-1}=\Lambda_{\pi(g)\cdot j}=\Lambda_{i}$.
+Thus $\bigoplus_{i\in I}\Lambda_{i}$ is in the subgroup generated by $g_{1},\ldots,g_{k}\cup\Lambda_{j}$.
+Now let $h\in\Gamma$.
+Note that there is $h^{\prime}\in G$ so that $\pi(h)=\pi(h^{\prime})$.
+Then $h^{-1}h^{\prime}\in\bigoplus_{i\in I}\Lambda_{i}$ completing the proof.
+∎
 Acknowledgments
+I would like to thank Francesco Fournier-Facio for many comments that shaped the Examples section.
+Most of the changes from the first version of this paper to the second came from a conversation I had with him.
+He explained a more elementary proof that Thompson’s group $F$ is not stable and suggested the other examples that appear in the paper.
 I would like to thank Eli Bashwinger for answering a question I had about the cohomology of Thompson and Thompson-like groups; his answer inspired me to make this note.